# ElastoGen: 4D Generative Elastodynamics

### v2

@@ 4.2 NeuralMTL & neural projection
which is an MLP whose weights can be pre-computed given the rasterized object. The output of $\mathcal{G}_{i}$ is then activated via a SVD module, which outputs $\mathbf{U}_{i}$, $\mathbf{V}_{i}$, and $\mathbf{S}_{i}$. As mentioned, $\mathbf{U}_{i}\mathbf{V}_{i}^{\top}$ constitutes the output of the local neural projection, but the energy check is performed through $\mathcal{N}$, which is embodied as a per-voxel compact convolution neural net. The weight coefficients of NeuralMTL are predicted by a generative diffusion model given the material type and parameters such as Young’s modulus $e$ and Poisson’s ratio $\nu$.
4.3 Decouple NeuralMTL from material parameters
As mentioned, NeuralMTL takes input as $\mathbf{F}_{i}$ and outputs $\mathcal{N}(\mathbf{F}_{i})$, a neural strain measure. This learned strain is then fit to equation 4 to check if ElastoGen reasonably minimizes equation 1 and is ready for the next time step. NeuralMTL is expected to fully accommodate material nonlinearity. Therefore, different material parameters $\{e,\nu\}$ guide NeuralMTL to yield different outputs even under the same $\mathbf{F}_{i}$. A straightforward approach is to train NeuralMTL $\mathcal{N}(\mathbf{F}_{i},e,\nu)$ directly on both $\mathbf{F}_{i}$ and $\{e,\nu\}$. However, as NeuralMTL needs to be evaluated more frequently under different $\mathbf{F}_{i}$ (during the deformation) after the material parameters are given, we decouple the influences of $\mathbf{F}_{i}$ and $\{e,\nu\}$ to keep the network even more compact. Inspired by Zhang et al. (2024a), we note that NeuralMTL $\mathcal{N}(\mathbf{F}_{i})$ can be generated using another diffusion network guided by $\{e,\nu\}$ such that $\mathbf{W}=\mathcal{D}(e,\nu)$, where $\mathbf{W}$ is the parameters of the network $\mathcal{N}(\mathbf{F}_{i})$ and $\mathcal{D}$ is another diffusion model.
@@ 4.3 Decouple NeuralMTL from material parameters
To train the model $\mathcal{D}$, we prepare a dataset of paired $\{e,\nu\}$ and $\mathbf{W}$. To this end, we first uniformly sample both $e$ and $\nu$ at fixed intervals and then establish a topological order, as shown in the left inset. A target elastic energy $\Psi(e,\nu)$ can be easily computed for each sampled $\{e,\nu\}$. $\mathbf{W}$ is then obtained via the following optimization:
$\mathbf{W}=\underset{\mathbf{W}}{\mathrm{argmin}}\left\|\log(\frac{\omega_{i}}{2}\|\mathbf{F}_{i}\cdot\mathcal{N}(\mathbf{W},\mathbf{F}_{i})-\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\|^{2}+1)-\log(\Psi+1)\right\|^{2},$ (9)
@@
where $\mathcal{N}({\mathbf{W}},\mathbf{F}_{i})$ suggests parameters of $\mathcal{N}$ are prescribed by $\mathbf{W}$. We use the logarithmic function $\log$ to strongly penalize the energy deviation under the same deformation and to ensure that the energy is always non-negative. Since the energy function changes smoothly with $\{e,\nu\}$, our pre-defined topological order of $\{e,\nu\}$ samples greatly eases the training. $\mathbf{W}$ can converge within only hundreds of gradient descent iterations when training uses the previous $\mathbf{W}$ for initialization. During inference, after $\mathcal{D}$ predicts $\mathbf{W}$, we apply a few extra iterations of gradient descent to fine-tune these weights, ensuring $\mathcal{N}$ fits the desired elastic energy function accurately. This two-step process ensures a smooth variation of the energy function with respect to $\{e,\nu\}$, allowing for efficient and precise generation of the network parameters.
4.4 Subspace encoding
@@
If the quadratic approximation of equation 1 is exact, NeuralMTL, $\mathcal{N}$, is not needed. After obtaining $\mathbf{Q}_{i}$ for all voxels, we set its derivative to zero leading to:
$\left(\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}\right)\mathbf{q}_{n+1}=\mathbf{f}_{q}+\frac{\mathbf{M}}{h^{2}}(\mathbf{q}_{n}+h\dot{\mathbf{q}}_{n})+\sum_{i}\mathbf{b}_{i},$ (10)
@@
where $\mathbf{b}_{i}=\mathbf{L}_{i}\mathbf{q}_{n}-\frac{\partial E_{i}}{\partial\mathbf{q}}$. We refer to $\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}$ as the global matrix, which is constant in this case. As a result, one can perform a pre-factorization converting the global matrix into lower and upper triangles to facilitate an effective solve of the linear system. However, the use of NeuralMTL alters the energy landscape nonlinearly, which makes $\mathbf{L}_{i}(\mathbf{q})$ dependent on the current deformed pose $\mathbf{q}$. Evaluating the system in a full implicit manner requires the information of $\nabla_{\mathbf{q}}\mathbf{L}_{i}$ and thus $\nabla_{\mathbf{q}}\mathcal{N}$, which is not only prohibitive but also less stable as extra training constraints need to be imposed e.g., to penalize $|\nabla\mathcal{N}|$ to prevent overfitting. To this end, we employ a lagged approach in computing $\mathbf{L}_{i}$ by using $\mathbf{q}$ from the most recent update.
Figure 2: Deforming object with the rasterization grid.
Solving the global matrix in a neural network way is challenging as all the features at vertices will become densely correlated via the matrix inverse, even the global matrix itself is sparse. To deal with this difficulty and to ensure EnasltoGen produces a physically accurate trajectory of the object, we decompose the effect of the global solve over $\mathbf{q}$ by applying multiple local operators at $\mathbf{q}_{i}$. Each local process works like a Laplacian operator, smoothing the rhs of equation 10 that depends on neural projection results of $\mathbf{P}_{i}$. Conceptually, this strategy can also be understood as finding a way to solve a linear system of the global matrix in a matrix-free manner.
Following this inspiration, we build ElastoGen as a two-level RNN network. The outer level of RNN or RNN-1 (e.g., see figure 1) repeats local NeuralMTL adjustments over $\mathcal{G}_{i}$ at each voxel region and the neural projection for $\mathbf{P}_{i}$. The inner RNN i.e., RNN-2 tackles the global smoothing. Specifically, we repeat a local smoothing conventional kernel and approximate the global smoothing effect as the outcome of repetitive local smoothing. Each local smoothing relaxes or releases the concentrated strain predicted by NeuralMTL $\mathcal{N}$ via expanding or shrinking its interface, which is shared with its neighboring voxels so that the local relaxation is slowly propagated over the entire object. Since a local relaxation is always applied at a voxel with eight vertices in our implementation, the corresponding network module shares the same structure for all the voxels.
A drawback of this strategy lies in the fact that it often takes a large number of RNN loops to generate a good global relaxation result. This is because local operations are more effective in processing locally concentrated strains, while object-wise global deformation can only be progressively approximated by information exchange via interface sharing across voxels. This is also a well-known limitation in numerical computation — Gauss-Seidel- or Jacobi-style iterative methods are less effective in relaxing low-frequency residual errors, which are often paired with a multigrid solver for large-scale problems.
We augment our RNN-2 with a deep encoder which extracts low-frequency strain the global matrix could generate. By encoding the input rhs of equation 10 into a low-dimension latent space of low-frequency deformations such as body-wise bending, twisting, or rotation, RNN-2 only needs to handle the remaining residual strains, which are often condensed locally. Determining the subspace encoding involves performing an SVD on the global matrix. Since our objects are rasterized, we use a rasterization grid as a general-purpose subspace. Each latent mode is visually similar to a gentle sine or cone wave e.g., see figure 2.
5 Experiments
We implement ElastoGen using Python. Specifically, we use PyTorch (Imambi et al., 2021) to implement the network and a simulator for training data generation. Our hardware platform is a desktop computer equipped with an Intel i7-12700F CPU and an NVIDIA 3090 GPU. Detailed statistics of the settings, models, and fitting errors are reported in table 1. All the experiments are also available in the supplemental video.
Table 1: Experiments statistics.  We report detailed settings of our experiments. ##\##DoFs: the average number of DOFs involved in the optimization. $\Delta t$: the size of timestep. ##\##R1: the average loop count of RNN-1 for each step. ##\##R2: the average number of RNN-2 loops for each timestep. # latent: the dimension of latent layer in the subspace encoder. EM: the elastic materials including Neo-Hookean (NH), StVK, and co-rotational (CR) models. Fitting error: the loss of NeuralMTL in equation 9. $t/\text{frame}$: the seconds needed for each frame.
Scene Grid resolution ##\##DoFs ##\##latent $\Delta t$#\##R1 ##\##R2 EM Fitting error $t/\text{frame}$
[…]
Cantilever  (Fig. 4) $16\times 3\times 3$ $432$ $18$ $0.001$ $5$ $108$ All $4.11\times 10^{-4}$ $0.01$
Cantilever (Fig. 8) $16\times 3\times 3$ $432$ $18$ $0.001$ $15$ $140$ NH $9.67\times 10^{-5}$ $0.01$
[…]
5.1 4D generation for any shapes
ElastoGen generates 4D elastic dynamics of 3D models with any shapes. To demonstrate this, we conduct experiments on multiple models from ShapeNet (Chang et al., 2015) with arbitrary external forces and boundary conditions. Some results of ElastoGen are shown in figure 3, and more are available in the appendix. All 3D objects are rasterized with a $32\times 32\times 32$ grid, which also serve as our subspace encoding. Cabinets are fixed at the bottom, twisted, and then released to yield elastic oscillations. Towers and plants sway under prescribed wind fields. Airplanes are pinned at the middle. Users apply sharp dragging force at the tip of the wings, resulting in interesting and realistic dynamic effects. These results show that different boundary conditions and external forces produce plausible dynamic outcomes.
Figure 3: ElastoGen on ShapeNet.  ElastoGen generates physically grounded 4D dynamics for objects of any geometries. To demonstrate this property, we run ElastoGen for a wide range of 3D objects in ShapeNet with different boundary conditions and external forces. This figure shows snapshots of a subset of our results including cabinets (green), towers (blue), plants (yellow), and airplanes (red). These experiments are under the rasterization resolution of $32\times 32\times 32$.
5.2 Quantitaive validation of NeuralMTL
ElastoGen replicates the behavior of real-world and complicated hyperelastic materials with different material parameters. We quantitatively compare the results generated with ElastoGen and simulated using the finite element method (FEM). We report the comparison for a standard bending test of a cantilever beam. We use ElastoGen to predict the further trajectory for three classic materials co-rotational (Brogan, 1986), Neo-Hookean (Wu et al., 2001), and StVK (Barbič & James, 2005). More general nonlinear materials, such as spline-based materials (Xu et al., 2015), are also supported. Each material is tested with three different Poisson’s ratios while keeping a fixed Young’s modulus (Poisson’s ratio alters the material response more nonlinearly than Young’s modulus). The results of ElastoGen, as shown in figure 4 (b), align well with the results obtained from the classic method of FEM. Both overlap nearly perfectly. Such superior accuracy is due to our NeuralMTL prediction. As shown in figure 4 (a), the diffusion-generated strain from NeuralMTL closely matches the ground truth (GT) with the correlation coefficient $r$ being larger than $0.98$ (calcualted as $r=\frac{\sum_{i=1}^{n}\left(g_{i}-\bar{g}\right)\left(f_{i}-\bar{f}\right)}{\sqrt{\sum_{i=1}^{n}\left(g_{i}-\bar{g}\right)^{2}\sum_{i=1}^{n}\left(f_{i}-\bar{f}\right)^{2}}}$ for each sample point $f_{i}$ and $g_{i}$ on neural strain and the ground truth curve, and $\bar{f}$ and $\bar{g}$ are their averages). We also plot the total neural energy variation over time for those materials ($\nu=0.32$) in figure 4 (c).
Figure 4: Quantitative validation of NeuralMTL.  (a) Comparison between the energy computed from NerualMTL strain and the ground truth energy. (b) Comparison with FEM under different material parameters. The relative positional error between ElastoGen (solid bars) and ground truth (red wireframes) is less than $5\%$. (c) Plots of the elastic energy during the prediction.
5.3 Versatility
ElastoGen is a general-purpose generative AI model. As long as a 3D object can be rasterized, ElastoGen deals with both explicit, e.g., as shown in figure 3, and implicit shape representations. For instance, when ElastoGen readily takes an implicit neural radiance field (NeRF) (Mildenhall et al., 2021) based model. One can conveniently employ the Poisson-disk sampling as described in Feng et al. (2023) to obtain the rasterized model. Given user-specified external forces or position constraints, ElastoGen generates its further dynamics directly via a neural network without resorting to an underlying physic simulator as used in PIE-NeRF (Feng et al., 2023). Similarly, a 3DGS (3D Gaussian splatting)-based model (Kerbl et al., 2023b) can also feed to ElastoGen for 4D generation. To show this, we report two experiments using multiple view images from the NeRF datasets as the input to ElastoGen in figure 5.
ElastoGen can benefit artists and animators by quickly producing high-quality 4D animations even for complicated models. We show such examples in figure 7 of two high-resolution objects discretized as triangle meshes. ElastoGen produces visually pleasing and physically accurate dynamics while preserving the dynamic details of the fine structures. Please refer to the supplementary video for more details. We can also inversely learn the material parameter from the video to make the generation consistent with the observation.
@@ 5.3 Versatility
Figure 5: ElastoGen with implicit models.  ElastoGen is compatible with both explicit and implicit models. We dense-sample the space of an implicit neural field to obtain its rasterization. Instead of running a physics simulator, ElastoGen directly yields physically accurate dynamics of the implicit model, which can be synthesized from novel camera poses. This enables a direct image-to-image generation.
5.4 More comparisons & ablation study
@@
Comparison with ground truth.   In addition to figure 4, we further compare ElastoGen with the FEM simulation under large-scale nonlinear twisting. The comparison is based on the Neo-Hookean material. For highly nonlinear instances, the physical accuracy of ElastoGen relies on the RNN loops — more loops at both RNN-1 and RNN-2 effectively converge ElstoGen to the ground truth. Nevertheless, for general-purpose generation, fewer iterations also yield good results. The detailed experiment and error plots are reported in figure 8.
@@
Figure 6: Comparison (trajectory) between ElastoGen, Gen-2 (Inc., ) and PhysDreamer (Zhang et al., 2024b).  We visualize the trajectory of a swinging carnation using ElastoGen, Gen-2, and PhysDreamer. Note that PhysDreamer can only produce plausible elastodynamics with tiny time steps ($\Delta t<6.0\times 10^{-5})$.
@@
Comparison with SOTA competitors.  We further compare ElastoGen with existing 4D generative models including Gen-2 (Inc., ) and PhysDreamer (Zhang et al., 2024b). ElastoGen demonstrates superior physical accuracy and geometric consistency. Specifically, Gen-2 produces a moderate movement with very little nonlinearity like rotation and bending. In contrast, ElastoGen successfully synthesizes physically accurate large-scale motion. Gen-2 fails to maintain geometric consistency over time. Both the color of the flower and the geometry of the stem have changed using Gen-2. This is a common issue for observation-based 4D generative models, where visual correlations in training data are highly complex and challenging to be decoupled by a monolithic deep model. Note that PhysDreamer can only produce plausible elastodynamics with tiny time steps ($\Delta t<6.0\times 10^{-5})$ due to the underlying explicit integration, which is known to be unstable under large time steps. In contrary, ElastoGen is able to generalize on large time steps. In table 5.4, we present a quantitative comparison of error using the Intersection over Union (IoU) metric between ElastoGen, Gen-2 (Inc., ), and PhysDreamer (Zhang et al., 2024b). The reference data is generated using Feng et al. (2023). Our method demonstrates superior accuracy in comparison to the others.
Figure 7: ElastoGen on complex scenes.  ElastoGen seamlessly accommodates complex meshes with intricate geometries and fine structures. With subspace encoding, ElastoGen preserves both high-frequency local details and low-frequency model-wise deformations.
Convergence study.  To quantify the impact of RNN loops and the subspace encoding on results, we compare ElastoGen predictions using different RNN loops with the ground truth, computed via solving the global matrix with a direct solver, in terms of relative error.
@@
[…]
tableComparison of quantative error between ElastoGen, Gen-2 and PhysDreamer.  We compute the Intersection over Union (IoU) using reference data generated by Feng et al. (2023). Higher IoU values indicate greater accuracy.
The results and convergence plots are shown in figure 8. In this standard test, one end of the beam is fixed, and ElastoGen predicts its twisting trajectory under external forces. We note that 50 RNN loops converge ElastoGen prediction to GT. Aggressively decreasing the loop count to 20 still yields satisfactory results. In contrast, 1, 3, and 5 iterations result in noticeably stiffer dynamics. In this experiment, RNN-2 uses an 18-dimension subspace encoder to extract low-frequency residuals. Without the encoding, local relaxation fails to converge.
6 Conclusion
ElastoGen is a knowledge-driven deep model that embeds physical principles and numerical procedures into the network design. As a result, EasltoGen is surprisingly lightweight and compact. Each module is tailored for a well-defined computational task for minimizing the total variational energy. This design allows for decoupled training, eliminating the need for large-scale training datasets. The accuracy of ElasoGen can be easily controlled by NeuralMTL which predicts the current strain from observed numerical computations.
ElastoGen also has limitations. The current version of ElastoGen lacks the support for collisions. It becomes less efficient for thin geometry as many convolution operations is applied on empty voxels. ElastoGen may fail to converge with extremely stiff materials like a near-rigid object. In the future, we plan to keep enhancing the scope of ElastoGen e.g., by integrating dynamics for more physical phenomena such as fluids, granular materials and plasticity, adding collision support, and automating the setting of physical parameters to ultimately achieve the goal of generating real-world dynamics with mouse clicks.
Appendix A Appendix
A.1 Supplemental video

### v3

@@ Introduction
which is a convolutional neural network (CNN) whose weights can be pre-computed given the rasterized object. The output of $\mathcal{G}_{i}$ is then activated via a SVD module, which outputs $\mathbf{U}_{i}$, $\mathbf{V}_{i}$, and $\mathbf{S}_{i}$.
And the NeuralMTL $\mathcal{N}$ can also be implemented as a CNN whose weights are predicted by a generative diffusion model given the material type and parameters such as Young’s modulus $e$ and Poisson’s ratio $\nu$.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 2: (a). NeuralMTL learns a mapping $\mathcal{N}$ to warp $\mathbf{F}_{i}$, enabling the quadratic strain energy to work for hyperelastic materials (b). Sampling topology over 2D input space to ensure smooth output variation. (c) Deforming object with rasterization grid.
Decoupling NeuralMTL from deformation and material parameters
NeuralMTL takes $\mathbf{F}_{i}$ as input and outputs the neural strain measure $\mathcal{N}(\mathbf{F}_{i})$. This strain is then evaluated against Eq. (4) to check if ElastoGen reasonably minimizes Eq. (1) and is ready for the next time step. NeuralMTL is designed to accommodate material nonlinearity, meaning different material parameters $\{e,\nu\}$ result in different outputs even for the same $\mathbf{F}_{i}$.
A direct approach would be training NeuralMTL $\mathcal{N}(\mathbf{F}_{i},e,\nu)$ on both $\mathbf{F}_{i}$ and $\{e,\nu\}$. However, since NeuralMTL needs to be evaluated frequently during deformation, we decouple the influences of $\mathbf{F}_{i}$ and $\{e,\nu\}$ to maintain a compact network. Following the approach in (zhang2024metadiff), we generate $\mathcal{N}(\mathbf{F}_{i})$ using a diffusion model $\mathcal{D}(e,\nu)$, where $\mathbf{W}=\mathcal{D}(e,\nu)$ is the parameters for the network $\mathcal{N}(\mathbf{F}_{i})$.
@@
To train the model $\mathcal{D}$, we prepare a dataset of paired $\{e,\nu\}$ and $\mathbf{W}$. To this end, we first uniformly sample both $e$ and $\nu$ at fixed intervals and then establish a topological order, as shown in Fig. 2 (b). A target elastic energy $\Psi(e,\nu)$ can be easily computed for each sampled $\{e,\nu\}$. $\mathbf{W}$ is then obtained via the following optimization:
$\mathbf{W}=\underset{\mathbf{W}}{\mathrm{argmin}}\left\|\log(\Psi_{\mathcal{N}}+1)-\log(\Psi+1)\right\|^{2},$ (9)
@@
where $\Psi_{\mathcal{N}}=\frac{\omega_{i}}{2}\|\mathbf{F}_{i}\cdot\mathcal{N}(\mathbf{W},\mathbf{F}_{i})-\mathbf{U}_{i}\mathbf{V}_{i}^{\top}\|^{2}$, and $\mathcal{N}({\mathbf{W}},\mathbf{F}_{i})$ suggests parameters of $\mathcal{N}$ are prescribed by $\mathbf{W}$. The logarithmic function $\log$ is used to penalize energy deviations under the same deformation, ensuring non-negative energy. As the energy function varies smoothly with $\{e,\nu\}$, our predefined topological ordering of $\{e,\nu\}$ samples facilitates efficient training. $\mathbf{W}$ can converge within a few hundred gradient descent iterations when initialized from the previous $\mathbf{W}$. During inference, after $\mathcal{D}$ predicts $\mathbf{W}$, we perform a few additional gradient descent iterations to fine-tune the weights, ensuring $\mathcal{N}$ accurately fits the desired elastic energy function. This two-step process enables smooth variation of the energy function with respect to $\{e,\nu\}$, ensuring efficient and precise network parameter generation.
Subspace encoding
@@
If the quadratic approximation of Eq. (1) is exact, NeuralMTL, $\mathcal{N}$, is not needed. And after obtaining $\mathbf{Q}_{i}$ for all voxels, we set its derivative to zero leading to:
$\left(\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}\right)\mathbf{q}^{n+1}=\mathbf{f}_{q}+\frac{\mathbf{M}}{h^{2}}(\mathbf{q}^{n}+h\dot{\mathbf{q}}^{n})+\sum_{i}\mathbf{b}_{i},$ (10)
@@
where $\mathbf{b}_{i}=\mathbf{L}_{i}\mathbf{q}_{n}-\frac{\partial E_{i}}{\partial\mathbf{q}}$. The term $\frac{\mathbf{M}}{h^{2}}+\sum_{i}\mathbf{L}_{i}$ as the global matrix, which remains constant in this case. This allows for pre-factorization, converting the global matrix into lower and upper triangular forms to facilitate efficient solving of the linear system. However, the use of NeuralMTL introduces nonlinearities in the energy landscape, making $\mathbf{L}_{i}(\mathbf{q})$ dependent on the current deformed pose $\mathbf{q}$. A fully implicit evaluation would require $\nabla_{\mathbf{q}}\mathbf{L}_{i}$ and $\nabla_{\mathbf{q}}\mathcal{N}$, which is computationally expensive and less stable due to the need for additional training constraints (e.g., penalizing $|\nabla\mathcal{N}|$ to prevent overfitting). To address this, we adopt a lagged approach, computing $\mathbf{L}_{i}$ by using $\mathbf{q}$ from the most recent update.
Directly inverting the global matrix is computationally expensive due to the dense correlations between vertices. As an alternative, traditional iterative methods update each node by incorporating information from its neighbors, using results from the previous iteration, and continue relaxing locally until convergence is achieved. Inspired by these methods, we decompose the global solve for $\mathbf{q}$ by perform multiple local operations at $\mathbf{q}_{i}$. Additionally, since the objects are rasterized into a grid, these repetitive local operators can be efficiently implemented using a recurrent CNN. This CNN takes $\mathbf{q}_{i}$ as input and outputs the relaxed result after one step. This CNN collects information from the vertices to the voxels via a convolution, then relaxes it back to the vertices using a transposed convolution. Additional implementation details can be found in our supplementary material. Conceptually, this approach functions as a matrix-free method for solving the global system, ensuring that ElastoGen generates a physically accurate trajectory.
ElastoGen is therefore built as a two-level RNN network. The outer RNN (RNN-1, see Fig. 1) applies local NeuralMTL adjustments over $\mathcal{G}_{i}$ at each voxel region and the neural projection for $\mathbf{P}_{i}$. The inner RNN (RNN-2) handles global solve by iterating a CNN, approximating the local relaxation effect through repeated local operations. Each local operator relaxes the concentrated strain predicted by NeuralMTL $\mathcal{N}$ and is propagated across the object. The process iterates until the difference between the results of two consecutive iterations is below a specified threshold.
@@
Figure 3: ElastoGen with implicit models.  ElastoGen supports both explicit and implicit models. We dense-samples the implicit neural field, directly generating physically accurate dynamics without a simulator, enabling image-to-image generation from novel camera poses.
However, this approach requires many RNN loops to achieve effective global relaxation. This is because local operations are more effective in processing locally concentrated strains, while object-wise global deformation can only be progressively approximated by information exchange across voxels. This is also a well-known limitation in numerical computation — Gauss-Seidel- or Jacobi-style iterative methods are less effective in relaxing low-frequency residual errors, which are often paired with a multigrid solver for large-scale problems.
To address this issue, we apply SVD to encode Eq. (10) into a low-frequency latent space and directly solve the system within this space. The result is then decoded back to the original space, serving as the initialization for RNN-2. However, since the global matrix varies across objects and timesteps, performing frequent SVD computations on the global matrix is inefficient. Given that only a prediction of the overall low-frequency deformation is required, we approximate the object’s global matrix using the global matrix of a uniform rasterized grid. This projection can be precomputed, eliminating the need for recalculation at each step or for each object. With this approach, each latent mode takes the form of a smooth sine wave, as illustrated in Fig. 2 (c).
Experiments
We implement ElastoGen using Python. Specifically, we use PyTorch (imambi2021pytorch) to implement the network and a simulator for training data generation. Our hardware is a desktop computer equipped with an Intel i7-12700F CPU and an NVIDIA 3090 GPU. Detailed statistics of the settings, models, and fitting errors are reported in Tab. 1. All the experiments are also available in the supplemental video.
Scene Grid resolution #\#DoFs #\#latent $\Delta t$ #\#R1 #\#R2 EM Fitting error $t/\text{frame}$
Cantilever  (Fig. 4)
$16\times 3\times 3$ $432$ $18$ $0.001$ 55 $108$ All $4.11\times 10^{-4}$ $0.01$
Lego (Fig. 3)
$26\times 46\times 30$ $11\text{K}$ $54$ $0.005$ $15$ $320$ NH $2.34\times 10^{-4}$ $0.44$
Drums (Fig. 3)
$28\times 22\times 34$ $4\text{K}$ $54$ $0.005$ $15$ $320$ CR $7.63\times 10^{-5}$ $0.21$
Bridge (Fig. 5)
$66\times 13\times 27$ $7\text{K}$ $81$ $0.003$ 55 $96$ StVK $5.78\times 10^{-4}$ $0.92$
Ship (Fig. 5)
$53\times 33\times 16$ $14\text{K}$ $81$ $0.001$ 55 $100$ NH $2.34\times 10^{-4}$ $1.20$
monster (Fig. 6)
$32\times 30\times 22$ $20\text{K}$ $36$ $0.001$ 55 $100$ NH $2.34\times 10^{-4}$ $0.78$
shoe (Fig. 6)
$48\times 30\times 20$ $28\text{K}$ $36$ $0.001$ 55 $100$ NH $2.34\times 10^{-4}$ $1.08$
Table 1: Experiments statistics.  We report detailed settings of our experiments. #\#DoFs: the average number of DOFs involved in the optimization. $\Delta t$: the size of timestep. #\#R1: the average loop count of RNN-1 for each step. #\#R2: the average number of RNN-2 loops for each timestep. # latent: the dimension of latent layer in the subspace encoder. EM: the elastic materials including Neo-Hookean (NH), StVK, and co-rotational (CR) models. Fitting error: the loss of NeuralMTL in Eq. (9). $t/\text{frame}$: the seconds needed for each frame.
Quantitative validation of NeuralMTL
ElastoGen replicates the behavior of complex hyperelastic materials with varying parameters. We quantitatively compare ElastoGen’s results with those from the finite element method (FEM) using a cantilever beam bending test. ElastoGen predicts the trajectories for three classic materials co-rotational (brogan1986corotation), Neo-Hookean (wu2001nh), and StVK (barbivc2005stvk). Each material is tested at three Poisson’s ratios, with a fixed Young’s modulus. (Poisson’s ratio alters the material response more nonlinearly than Young’s modulus). The results of ElastoGen, as shown in Fig. 4 (b), align well with the results obtained from the classic method of FEM. Both overlap nearly perfectly. Such superior accuracy is due to our NeuralMTL prediction. As shown in Fig. 4 (a), the diffusion-generated strain from NeuralMTL closely matches the ground truth (GT) with the correlation coefficient $r$ being larger than $0.98$ (calculated as $r=\frac{\sum_{i=1}^{n}\left(g_{i}-\bar{g}\right)\left(f_{i}-\bar{f}\right)}{\sqrt{\sum_{i=1}^{n}\left(g_{i}-\bar{g}\right)^{2}\sum_{i=1}^{n}\left(f_{i}-\bar{f}\right)^{2}}}$ for each sample point $f_{i}$ and $g_{i}$ on neural strain and the ground truth curve, and $\bar{f}$ and $\bar{g}$ are their averages). We also plot the total neural energy variation over time for those materials ($\nu=0.32$) in Fig. 4 (c).
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 4: Quantitative validation of NeuralMTL.  (a) Comparison between the energy computed from NeuralMTL strain and the ground truth energy. (b) Visualization of the final static state of the cantilever beam generated by our method, with different materials and material parameters. (c) Plots of the elastic energy during the prediction.
Figure 5: Experiments on high-resolution mesh objects.  ElastoGen generates realistic nonlinear elastic dynamics under external forces. Both high-frequency local details and low-frequency global deformations are effectively captured with a carefully designed nested RNN architecture.
Versatility
ElastoGen is a general-purpose generative AI model. As long as a 3D object can be rasterized, ElastoGen deals with both explicit, e.g., as shown in Fig. 5, and implicit shape representations. For instance, when ElastoGen readily takes an implicit neural radiance field (NeRF) (mildenhall2021nerf) based model. One can conveniently employ the Poisson-disk sampling as described in feng2023pie to obtain the rasterized model. Given user-specified external forces or position constraints, ElastoGen generates its further dynamics directly via a neural network without resorting to an underlying physic simulator as used in PIE-NeRF (feng2023pie). Similarly, a 3DGS (3D Gaussian splatting)-based model (kerbl20233d) can also feed to ElastoGen for 4D generation. We demonstrate this using NeRF dataset images in Fig. 3.
ElastoGen enables artists and animators to quickly produce high-quality 4D animations, even for complex models. Examples of high-resolution, triangle mesh objects are shown in Fig. 5, where ElastoGen generates visually accurate dynamics while preserving fine structural details. Additionally, material parameters can be inversely learned from video to ensure consistency with observed dynamics.
ElastoGen can also serve as a downstream model for 3D generation frameworks to accomplish true 4D generation tasks. After rasterizing the 3D object generated by the upstream model, it can be used as the input to ElastoGen, which then produces physically plausible animations driven by forces or constraints. As shown in Fig. 6, by integrating ElastoGen with ARM (feng2024arm), we achieve high-quality and physically accurate 4D elastodynamics generation.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: 4D generation with ARM (feng2024arm).  We combine ElastoGen with a 3D generation model ARM (feng2024arm) to perform 4D elastodynamics generation. The first row shows a monster shaking its head, while the second row shows a shoe being squeezed.
More comparisons
@@
Comparison with ground truth. In addition to Fig. 4, we further compare ElastoGen with the FEM simulation under large-scale nonlinear twisting. The comparison is based on the Neo-Hookean material. For highly nonlinear instances, the physical accuracy of ElastoGen relies on the RNN loops — more loops at both RNN-1 and RNN-2 effectively converge ElastoGen to the ground truth. Nevertheless, for general-purpose generation, fewer iterations also yield good results. Detailed experimental results and error plots are provided in the supplementary materials.
Comparison with SOTA competitors. We compare ElastoGen with existing 4D generative models, including Gen-2 (runwayml2024image2video_gen2) and PhysDreamer (zhang2024physdreamer) in Fig. 7, showing that ElastoGen excels in both physical accuracy and geometric consistency. Gen-2 generates moderate motion with limited nonlinearity, such as rotation and bending, and fails to maintain geometric consistency over time,
@@
Figure 7: Comparison (trajectory) between ElastoGen, Gen-2 (runwayml2024image2video_gen2) and PhysDreamer (zhang2024physdreamer).  We visualize the trajectory of a swinging carnation using ElastoGen, Gen-2, and PhysDreamer. Note that PhysDreamer can only produce plausible elastodynamics with tiny time steps.
@@
causing changes in both color and shape. This is a common issue for observation-based 4D models, where complex visual correlations in training data are difficult to decouple in a monolithic deep model. PhysDreamer produces plausible elastodynamics only at small time steps ($\Delta t<6.0\times 10^{-5})$ due to the underlying explicit integration, which is known to be unstable under large time steps. In contrary, ElastoGen is able to generalize on large time steps. In Tab. 2, we present a quantitative comparison of error using the Intersection over Union (IoU) metric between ElastoGen, Gen-2 (runwayml2024image2video_gen2), and PhysDreamer (zhang2024physdreamer). The reference data is generated using (feng2023pie). Our method demonstrates superior accuracy in comparison to the others.
@@
Table 2: Comparison of quantitative error between ElastoGen, Gen-2 and PhysDreamer.  We compute the Intersection over Union (IoU) using reference data generated by (feng2023pie). Higher IoU values indicate greater accuracy.
Conclusion
ElastoGen is a knowledge-driven deep model that embeds physical principles and numerical procedures into the network design. As a result, it is remarkably lightweight and compact. Each module is designed for a specific computational task aimed at minimizing the total variational energy. This modular design enables decoupled training, removing the need for large-scale training datasets. The accuracy of ElastoGen can be easily controlled by NeuralMTL, which predicts the current strain based on observed numerical computations.
ElastoGen also has limitations. It currently lacks collision support and is less efficient for thin geometries due to excessive convolution operations on empty voxels. It may also fail to converge with highly stiff materials, such as near-rigid objects. Future improvements will include adding dynamics for more physical phenomena, integrating collision support, and automating the setting of physical parameters to enable real-world dynamics with minimal input.
Appendix A Supplemental video